The veocity v(t) of a particle as a function of time is given by v(t)= (2.3m/s)+(4.1m/s^2)t-(6.2m/s^3)t^2. What is the average acceleration of the particle between t=1.0s and t=2.0s?
Answer #1

VW) = 2.3+ 4.1t - 6.27 v(t2) - VCH) Average acceleration Covers tz-t, [2-344-1x 2 -6.2x2]- (2-3X4.1X1–62x1} = 2-1 4.1X1 -6.2